# Interactive Agents to Overcome Ambiguity in Software Engineering

### v1

@@ 3.2 Leveraging Interaction in Ambiguity
Using interaction, the Claude Sonnet and Haiku agents recreate 80% of the performance in the Full setting. However, with Deepseek and Llama 3.1, the relative performance is lower, of 59% and 54%, respectively. Claude Sonnet 3.5’s high resolve rate in the Hidden setting is likely due to its superior programming acumen. The performance is surprising, as a human would be able to decipher little about the expectations given the summarized issue. Better programming models can potentially extract more information from the stack trace by reproducing the error themselves. We observe that the Claude Haiku model achieves a performance relative to the Full setting similar to that of the Claude Sonnet model, despite having inferior coding abilities. Thus, there is no direct correlation between the number of parameters or coding ability and a model’s ability to leverage interaction. This hints towards better training practices that can lead to better integration of the new information.
@@
Table 1: Wilcoxon Signed-Rank Test Results for Hidden vs Interaction setting and for Interaction vs Full setting for each model
This experiment highlights the importance of interaction in mitigating ambiguity. Since many real-world software engineering problems are underspecified (Chowdhury et al., 2024), interactive systems are essential for ensuring alignment and reducing safety risks. However, current models default to non-interactive behavior even when faced with ambiguity and struggle to match the performance seen in well-specified settings. While interactive trajectories show performance gains over non-interactive approaches for ambiguous inputs, the improvement is not statistically significant, indicating strong potential for improvement.
3.3 Impact of Interaction Details on Model Performance
Model Navigational Info (%) Resolve Rate With (%) Resolve Rate Without (%)
Claude Sonnet 3.5 8.96 59.52 37.94
Claude Haiku 3.5 24.67 36.94 24.78
Deepseek-v2 30.70 13.19 4.62
Llama 3.1 70B 30.28 6.34 4.28
Table 2: The percentage of issues where navigational information was acquired in the Interaction setting, along with the corresponding resolve rates with and without navigational information. Navigational information refers to file locations requiring modification, helping to avoid tedious code exploration. The resolve rates offer insight into how the information obtained during interaction impacts overall performance.
In the Interaction setting of the previous experiment, the information gained can be broadly categorized into two types: informational, which relates to the expected behavior or nature of the error, and navigational, which pertains to the locations of the files to modify. While informational details are typically obtained in nearly every interaction, the models request navigational details less frequently. We measure the resolve rates separately for instances where the model asks for navigational details and when it does not, examining the impact on performance when models must rely only on informational details versus when navigational details are also accessible.
As seen in Table 2, requesting navigational details improves performance across all models by providing cues beyond described behavior and errors. However, some models rely too heavily on this information and struggle when it’s missing. Smaller models like Llama 3.1 and Deepseek-v2 request file locations more often but underperform without them. Claude models, particularly Sonnet, better leverage informational cues, achieving higher resolve rates even without navigational details. Deepseek, by contrast, performs worse than its Hidden setting when file locations are absent, highlighting its dependence. This reliance leads to wasted turns searching for errors instead of identifying them efficiently. Llama 3.1 performs better than Hidden without file locations but gains little when they are provided, likely due to poor detail extraction (Section §4). Ideally, LLMs should generalize across diverse interaction types, as users may not always provide specific details, improving robustness in real-world software engineering tasks.
Takeaway: Interaction has significant potential to improve model performance in ambiguous tasks, but models, particularly the less strong open-weight models, struggle to leverage it effectively. Proprietary models like Claude Sonnet 3.5 and Haiku 3.5 achieve nearly 80% of their Full setting performance, with Haiku improving by 74% over its Hidden setting performance through effective integration of both informational and navigational cues. The lack of correlation between model size and its ability to utilize interaction suggests that better training practices play a more crucial role. In contrast, models like Deepseek-v2 and Llama 3.1 show limited gains, primarily due to their challenges in utilizing broader informational cues, which hinders their adaptability in ambiguous tasks.
4 RQ2: Ambiguity Detection
In real-world LLM and agent applications, task descriptions and prompts can vary in quality (Chowdhury et al., 2024). To detect ambiguity, a model must recognize unclear expectations or identify missing key information in its planned approach. However, interacting unnecessarily when sufficient information is already available can introduce inefficiencies and place an undue burden on the user. Here, we examine the capabilities of LLMs to detect ambiguous instructions in software engineering contexts.
4.1 Experimental Setup
In the instructions which outline the task, we present the agent with an option to interact during its solution trajectory and design three instructions with varying levels of encouragement to interact with the user. We track the input type the model chooses to interact with. The instructions, listed in order of increasing encouragement to interact, are: Neutral, where the agent is told it can ask questions if anything is unclear), Moderate Encouragement, where the agent is told to carefully check that all necessary information is available and only proceed after everything is clear, and Strong Encouragement, where the agent is told that asking questions is critical to task success (full prompts in §A).
Model Neutral Moderate Encouragement Strong Encouragement
Accuracy ↑ FPR ↓ FNR ↓ Accuracy ↑ FPR ↓ FNR ↓ Accuracy ↑ FPR ↓ FNR ↓
Claude Sonnet 3.5 0.60 0.00 0.81 0.84 0.24 0.09 0.76 0.36 0.10
Claude Haiku 3.5 0.54 0.00 0.97 0.57 0.02 0.90 0.63 0.06 0.66
Deepseek-v2 0.69 0.30 0.31 0.57 0.08 0.83 0.51 0.04 0.94
Llama 3.1 70B 0.48 0.46 0.57 0.47 0.95 0.09 0.52 0.93 0.06
Table 3: Model performance in ambiguity detection across prompts with varying levels of interaction encouragement. FPR refers to cases where the model interacted unnecessarily, while FNR refers to cases where it failed to interact when needed. A model that reliably distinguishes between underspecified and well-specified issues should have high accuracy, low FPR, and low FNR.
4.2 Effect of Different Prompts
Experiments to detect ambiguity demonstrate that, using prompt engineering, we can control the level of interaction with the user, as shown in Table 3. But this interactivity is not possible without clearly specifying it in the prompt wherein without any specific mention of interaction, the models almost never interact for any of the summarized issue inputs.
The Claude Sonnet model performs best with Moderate Encouragement, achieving the highest overall accuracy of 84% across all variations. Its counterpart from the same model family, Claude Haiku, is hesitant to interact even with Strong Encouragement. The Claude models show a drop in accuracy in cases where interaction is not needed as their overall interaction increases, indicating that the interaction fails to target underspecified inputs effectively. For the Deepseek model, we observe that the Neutral prompt gives the best results as interactivity surprisingly decreases with more encouragement. The accuracy in both the cases where interaction was desired and not desired is around 70%, which shows that the model is capable of distinguishing between well-specified and underspecified issues to some extent. The Llama model displays a greater, but arbitrary, tendency to interact across all prompts than other models.
4.3 Detection across Models
While interaction levels can be adjusted with prompting, both summarized issues and full issues have equal probability of being selected for interaction as interactivity increases, particularly with smaller models. Despite the stark difference in the language and detail of summarized issues and fully specified issues, the models, except Claude Sonnet, fail to reliably distinguish them, indicating that LLMs struggle to detect ambiguity even in obvious cases. All models, including Claude Sonnet, show big changes in the ambiguity detection behavior with prompt variations. Interestingly, Sonnet outperforms Haiku, likely due to its more extensive instruction tuning or Human Feedback training, which helps it better follow instructions and achieve the desired interactive trajectory. Surprisingly, even Deepseek adapts better to the task than Haiku.
Takeaway: Prompt engineering can influence model interactivity but fails to consistently improve ambiguity detection across models. When interaction is not explicitly prompted, models default to non-interactive behavior. Claude Sonnet shows some ability to distinguish ambiguous inputs, but other models, including Claude Haiku and Llama 3.1, struggle even with clear cues. This inconsistency reveals that models are not inherently equipped to detect underspecified tasks. Improving ambiguity detection requires dedicated training, not just prompt modifications.
5 RQ3: Question Quality
Figure 4: Agent questions and user responses to the same underspecified input are shown for Llama 3.1 70B, Deepseek-v2, and Claude Haiku 3.5. The examples highlight specific interaction patterns and differences in handling ambiguity. The corresponding model inputs are detailed in Table 5.
To gather missing information from underspecified inputs, the quality of an agent’s questions is crucial. While §3 evaluates task completion, the model performance in the experiment is influenced by the coding ability. Here, we focus solely on the quality of the questions posed by the agent to the user, measuring how effectively models extract relevant information under the assumption that users have the necessary details.
5.1 Experimental Setup
In this experiment, we evaluate the quality of the interactions between the agent and the user in the Interaction setting. We measure the novelty and detail level of the information obtained from the user’s answers to evaluate the quality, quantifying the new knowledge relative to the existing understanding of the agent. We employ two techniques to quantify the information obtained.
  1. 1.
    Cosine Distance: We compute the cosine distance ($1-cos(P,Q)$) between the embeddings of the summarized task $E_{\text{before}}$ and the cumulative knowledge after interaction with the user $E_{\text{after}}$ using a text embedding model. Lower distances indicate redundant user input, while higher values show meaningful information gain. We use OpenAI’s text-embedding-3-small as our embedding model.
  2. 2.
    LLM-as-judge (GPT-4o): The model scores the user answers on a scale of 1 to 5, where a higher score corresponds to more new and detailed information in the user’s response, such as specific files causing errors or function behavior. The prompt to the model includes the summarized issue, agent questions, and user responses for better context.
5.2 Information Gain from Interaction
For the quantitative evaluation of the quality of the question, both the cosine distance and the LLM-as-judge methods suggest a similar result, in which the Llama model performs significantly worse than the other models, whereas the other models achieve very similar information gains, as seen in Figure 5.
The Llama model has an average cosine distance of 0.101 when the embedding of the summarized issue is compared to the embedding of the user response appended to the summarized issue. Deepseek achieves the highest cosine distance of 0.142, while the Claude Sonnet and Haiku models achieve very similar cosine distances of 0.136 and 0.135.
Using LLM as a judge, we evaluate the specificity of the details present in the answers. Here again, the Llama 3.1 model achieves a significantly worse average score of 3.58 than the other models which see similar performance of around 4 out of 5.
Figure 5: Information Gain measured using (a) Cosine Distance Scores and (b) LLM-as-Judge Scores
5.3 Qualitative Analysis of Questions
The quantitative results can be further supported by a qualitative evaluation of the questions. Sample question-answer pairs reflecting common trends are shown in Figure 4. The Llama model asks fewer questions on average than other models in one message for user interaction, as seen in Table 4, and often poses overly general questions like, Are there any existing workarounds or temporary fixes?. These template-like questions are unproductive and less likely to gather useful information.
Deepseek, on the other hand, asks the most questions per message, allowing it to extract more information. Its questions, such as Are there any existing tests or examples that demonstrate the issue?, aim to extract, edge cases, documentation, or tests, and while common across multiple issues, they are reasonable and yield valuable details. But most questions are very specific and detailed, querying about the expected behavior. Often, due to the specificity of the question, the user might not have the required information.
Claude Sonnet asks fewer questions than Deepseek, likely because it explores the codebase first. The questions do not have easily discernible patterns and match the Deepseek model in specificity. The Haiku model, in contrast, follows a consistent template, typically asking three questions regardless of the input, although sub-questions may be present. Haiku’s questions are more keyword-driven based on the input, while Sonnet’s are based on a deeper understanding of the issue and codebase.
Model Avg. Number of Questions

### v2

@@ 3.2 Leveraging interaction in underspecificity
Using interaction, the Claude Sonnet 3.5 models and Haiku 3.5 recover up to 80% of the performance in the Full setting. However, with Deepseek, and Llama 3.1, the relative performance is lower, of 59%, and 54%, respectively. Claude Sonnet 4’s relative performance (61%) is lower than that of its predecessor, and absolute performance with interaction is also similar. Some models achieve higher resolve rates in the Hidden setting likely due to their superior programming acumen, or data leakage. Better programming models can potentially extract more information from the stack trace by reproducing the error themselves. Claude Sonnet 4 extensively explores the codebase and attempts multiple solutions to overcome the lack of information in the Hidden setting. On the other hand, Qwen 3 Coder displays unique behavior in this setting and relies on its internal knowledge for key insights about missing information (example in §A.7). These correct assumptions might inflate its performance in this setting. We observe that the Claude Haiku model achieves a performance relative to the Full setting similar to that of Claude Sonnet 3.5, despite having inferior coding abilities. Thus, there does not seem to be a direct correlation between the number of parameters or coding ability and a model’s ability to leverage interaction. This hints towards better training practices that can lead to better integration of the new information.
This experiment highlights the importance of interaction in handling underspecificity. Since many real-world software engineering problems are underspecified, interactive systems are essential for ensuring alignment and reducing safety risks. However, current models default to non-interactive behavior even when faced with severe lack of information and struggle to match the performance seen in well-specified settings. While interactive trajectories show performance gains over non-interactive approaches for underspecified inputs, there is a wide gap to the full performance, indicating strong potential for improvement.
3.3 Impact of interaction details on model performance
To better understand these differences in interaction effectiveness, we next examine what types of information models request and how they utilize it. In the Interaction setting of the previous experiment, the information gained can be broadly categorized into two types: informational, which relates to the expected behavior or nature of the error, and navigational, which pertains to the locations of the files to modify. The latter can be considered redundant information as it can be recovered from the codebase. While informational details are typically obtained in nearly every interaction, the models request navigational details less frequently. We measure the resolve rates separately for instances where the model asks for navigational details and when it does not, examining the impact on performance when models must rely only on informational details versus when navigational details are also accessible.
Model Nav Info (%) Resolve w/o Info (%) Resolve w/ Info (%)
Claude Sonnet 4 27.75 43.75 46.55
Qwen 3 Coder 18.60 47.60 40.20
Claude Sonnet 3.5 8.96 37.94 59.52
Claude Haiku 3.5 24.67 24.78 36.94
Deepseek-v2 30.70 4.62 13.19
Llama 3.1 70B 30.28 4.28 6.34
Table 1: % of issues where navigational information was acquired in the Interaction setting, and the resolve rates with and without it. Navigational information refers to file paths needing modification.
As seen in Table 1, requesting navigational details improves performance across most models by providing cues beyond described behavior and errors. However, some models rely too heavily on this information and struggle when it is missing. Smaller models like Llama 3.1 and Deepseek-v2 request file locations more often but underperform without them. With improvements in code localization ability, recent models like Claude Sonnet 4 and Qwen 3 Coder show lower performance boosts with this information. Qwen 3 Coder displays a unique behavior where its performance worsens after receiving file locations. An analysis of the trajectories reveals rigid behavior where the model gets the information from the user, yet proceeds to re-explore the code and come across the same information by itself, seemingly following a set protocol of approaching SWE-Bench style issues. This suggests that while the model acknowledges the user input, it does not easily modify its behavior, also evidenced by its need for stronger prompting to interact (§A.7). This rigid behavior wastes interaction turns on redundant navigational information, preventing the model from asking more valuable clarifying questions about task requirements. Claude models, particularly Sonnet 3.5, better leverage informational cues, achieving higher resolve rates even without navigational details. Deepseek, by contrast, performs worse than its Hidden setting when file locations are absent, highlighting its dependence. This reliance leads to wasted turns searching for errors instead of identifying them efficiently. Llama 3.1 performs better than Hidden without file locations but gains little when they are provided, likely due to poor detail extraction (Section §4). Ideally, LLMs should generalize across diverse interaction types, as users may not always provide specific details, improving robustness in real-world software engineering tasks.
Takeaway: While proprietary models like Claude Sonnet 3.5 and Haiku 3.5 effectively leverage interaction (recovering  80% of the performance gap), recent capable models show a shift: Claude Sonnet 4’s relative gains diminish despite stronger absolute performance, and Qwen 3 Coder rigidly adheres to predetermined protocols even when users provide explicit guidance. These patterns suggest that as models grow more capable, current training practices may inadequately prepare them to dynamically integrate interactive information, highlighting the need for approaches that prioritize adaptive behavior over task completion alone.
4 RQ2: Detection of incomplete task specifications
Can LLMs identify whether a given task description is missing crucial information? In real-world LLM and agent applications, task descriptions and prompts often vary in quality. Unnecessary interaction when sufficient information is already available can introduce inefficiencies and burden users. In this work, we evaluate whether LLMs can detect missing information in software engineering contexts by randomly presenting either fully-specified or underspecified issues, along with varying interaction prompts, and analyzing their interaction behavior across these conditions.
4.1 Experimental setup
In the instructions which outline the task, we present the agent with an option to interact during its solution trajectory and design three instructions with varying levels of encouragement to interact with the user. We track the input type the model chooses to interact with. The instructions, listed in order of increasing encouragement to interact, are: Neutral, where the agent is told it can ask questions if anything is unclear), Moderate Encouragement, where the agent is told to carefully check that all necessary information is available and only proceed after everything is clear, and Strong Encouragement, where the agent is told that asking questions is critical to task success (full prompts in Appendix §A).
Table 2: Model performance in underspecificity detection across prompts with increasing interaction encouragement. FPR: false positive rate (unnecessary interaction); FNR: false negative rate (missed necessary interaction). Ideal models have high accuracy, low FPR, and low FNR.
Model Neutral Moderate Strong
Acc ↑ FPR ↓ FNR ↓ Acc ↑ FPR ↓ FNR ↓ Acc ↑ FPR ↓ FNR ↓
Claude Sonnet 4 0.74 0.08 0.44 0.74 0.10 0.42 0.89 0.03 0.18
Qwen 3 Coder 0.50 0.00 1.00 0.50 0.00 1.00 0.50 0.00 1.00
Claude Sonnet 3.5 0.60 0.00 0.81 0.84 0.24 0.09 0.76 0.36 0.10
Claude Haiku 3.5 0.54 0.00 0.97 0.57 0.02 0.90 0.63 0.06 0.66
Deepseek-v2 0.69 0.30 0.31 0.57 0.08 0.83 0.51 0.04 0.94
Llama 3.1 70B 0.48 0.46 0.57 0.47 0.95 0.09 0.52 0.93 0.06
4.2 Effect of different prompts
Without explicit prompting, models almost never interact, even for severely underspecified inputs. Table 2 shows that prompt engineering can modulate interaction levels, but with highly variable effectiveness across models.
Claude family: Claude Sonnet 4 achieves the best performance with Strong Encouragement, representing substantial improvement over other models. Claude Sonnet 3.5 performs best with Moderate Encouragement (84% accuracy), while Claude Haiku 3.5 remains hesitant to interact even with strong prompting. The Sonnet models’ superior performance likely stems from better instruction-following capabilities.
Open-weight models show divergent behaviors: Deepseek-v2 exhibits counterintuitive behavior, performing best with Neutral prompting and degrading with stronger encouragement. Llama 3.1 shows excessive interaction across all prompts, interacting arbitrarily rather than strategically. Most critically, Qwen 3 Coder completely fails to interact under any condition (100% FNR across all prompts), achieving only chance-level accuracy (50%).
4.3 Detection across models
While interaction levels can often be adjusted with prompting, both summarized and full issues have equal probability of being selected for interaction as interactivity increases, particularly with smaller models. It is a concerning finding that even with stronger encouragement one of the leading open-weight models, Qwen 3, is rigid in its non-interactivity. Despite the stark difference in the language and detail of summarized issues and fully specified issues, the models, except Claude Sonnet, fail to reliably distinguish them, indicating that LLMs struggle to detect missing information even in obvious cases. Most models, including Claude Sonnet, show big changes in the detection behavior with prompt variations. Interestingly, the Sonnet models outperforms Haiku, likely due to superior instruction following capability, which helps it achieve the desired interactive trajectory. Surprisingly, Deepseek adapts better to the task than Haiku as well as Qwen 3.
Takeaway: Models default to non-interactive behavior unless explicitly prompted, yet prompt engineering alone proves insufficient for reliable underspecificity detection. While Claude Sonnet 4 demonstrates partial success, most models struggle to distinguish well-specified from underspecified tasks. Particularly concerning is Qwen 3 Coder’s complete non-responsiveness to interaction prompts (100% FNR), suggesting fundamental limitations in certain training approaches. These findings indicate that handling underspecified tasks requires dedicated training than prompt engineering alone.
5 RQ3: Question quality
Figure 4: Agent questions and user responses to the same underspecified input are shown for Llama 3.1 70B, Deepseek-v2, and Claude Haiku 3.5. They highlight specific interaction patterns and differences in handling missing information. The corresponding model inputs are detailed in Table 7.
Can LLMs generate meaningful and targeted clarification questions that gather the necessary information for task completion? To gather missing information from underspecified inputs, the quality of an agent’s questions is crucial. While §3 evaluates task completion, here we focus on how effectively models extract relevant information through their questions.
5.1 Experimental setup
We evaluate interaction quality in the Interaction setting using two complementary techniques:
  1. 1.
    Cosine distance: We compute the distance ($1-\cos(P,Q)$) between embeddings of the summarized task ($E_{\text{before}}$) and cumulative knowledge after interaction ($E_{\text{after}}$) using OpenAI’s text-embedding-3-small. Higher values indicate greater information gain.
  2. 2.
    LLM-as-judge (GPT-4o): Scores user answers on a 1-5 scale based on specificity and novelty of information (e.g., specific files, function behavior).
5.2 Information gain and question efficiency
Figure 5: Information Gain measured using Cosine Distance
Both metrics reveal that Llama 3.1 significantly underperforms (0.101 cosine distance, 3.58/5 LLM-judge score) compared to other models (Figure 5). More interesting are the patterns among stronger models: Qwen 3 Coder achieves the highest information extraction (0.179) but requires 50% more questions than Claude Sonnet 4 (6.02 vs 4.03, Table 6), yet both achieve similar resolve rates (46% vs 41.8%, Figure 3). Similarly, Claude Sonnet 3.5 and Haiku extract nearly identical information (0.136 vs 0.135) despite vastly different task performance (39.6% vs 26.8%). These disconnects reveal that how models integrate information matters as much as how much they extract.
Figure 6: Information Gain measured using LLM-as-Judge
The LLM-as-judge scores converge around 4/5 for all capable models (Figure 6), indicating they can elicit relevant information when prompted. However, cosine distance’s granularity reveals efficiency differences: similar information can be obtained with vastly different question quantities and strategies.
5.3 Question-asking strategies
Qualitative analysis of question-answer pairs (Figure 4) reveals three distinct approaches with different tradeoffs:
(1) Question quantity and user burden. Llama asks too few questions (2.61 avg.) with overly general phrasing ("Are there any existing workarounds?"), yielding minimal information. Qwen 3 asks the most (6.02 avg.), extracting maximum information but risking user overwhelm. As Table 1 shows, this high volume does not translate to better performance. Qwen’s resolve rate actually worsens with navigational information, suggesting rigid protocol-following rather than adaptive integration.
(2) Exploration efficiency. Claude Sonnet models (3.80-4.03 questions) achieve information gain comparable to Deepseek and Qwen (4.57-6.02 questions) by exploring the codebase first, then asking only what cannot be independently discovered. This exploration-first strategy produces nearly identical questions across Claude Sonnet 3.5 and 4 for some issues (Table 7), indicating consistent training. In contrast, Deepseek and Qwen ask immediately, including questions about implementation details recoverable from code.
(3) Answerability and specificity. Deepseek’s highly specific implementation questions often exceed user knowledge, wasting interaction turns. Claude targets behavioral aspects and concrete failure modes instead, better matching realistic user knowledge. Haiku follows a rigid three-question template regardless of context, while Sonnet adapts questions based on deeper issue understanding.
Takeaway: Effective clarification balances question quantity (avoiding user overwhelm), exploration efficiency (discovering what can be inferred before asking), and answerability (matching specificity to user knowledge). Claude Sonnet 4 achieves comparable information gain to Qwen (0.171 vs 0.179) with 50% fewer questions through exploration-first strategies, demonstrating that question quality and integration matter more than extraction volume.
6 Related Work
Code generation benchmarks
Ambiguity is a closely related domain to underspecificity, where model misinterpretation of user intent is a common failure mode. In both cases, clarification becomes necessary, though the causes differ. Ambiguity stems from vague or multi-interpretable inputs, while underspecificity arises when key information is entirely omitted. This is especially relevant in our setting, where models operate over intent summaries that may only partially capture user goals. Clarifying questions help mitigate ambiguity (Mu et al., 2023), and interactive, test-driven workflows generate test cases aligned with expectations, which users validate before code generation (Lahiri et al., 2023). Extensions of this approach employ runtime techniques to generate, mutate, and rank candidates based on user feedback (Fakhoury et al., 2024). Although effective, these workflows can burden users, highlighting the need to minimize intervention to essential cases.
Interactive ML systems
In interactive systems, ambiguity is often categorized and addressed via targeted clarification. Niwa and Iso (2024) introduces a taxonomy of instruction ambiguities, such as unclear output formats or contextual constraints, and applies disambiguation strategies accordingly. Similarly, Wang et al. (2024a) evaluates LLM behavior on ambiguous tool-use instructions, and Feng et al. (2024) uses reinforcement learning to optimize intervention. Although these systems successfully reduce ambiguity, underspecificity poses a subtler challenge, where there is missing context, leading to hallucinated assumptions and requires agents to clarify.
LLMs and ambiguity
Modern LLMs are not explicitly trained to resolve ambiguity via interaction (Zhang et al., 2024), but instruction tuning improves their performance when guided by prompt engineering (White et al., 2023). Ambiguity detection has been approached through uncertainty estimation (Zhang and Choi, 2023; Park et al., 2024) and self-disambiguation (Keluskar et al., 2024; Sterner, 2022; Sumanathilaka et al., 2024). For example, Kim et al. (2024) quantifies ambiguity using information gain. Although inference-only methods are cost-effective, they are less robust than training-based approaches for handling ambiguity. Chen et al. (2025) address disambiguation in conversational settings, but typically with only a single missing detail. In contrast, we study underspecification in complex agentic tasks, where multiple interdependent gaps can arise dynamically, and agents may take many steps before recognizing missing information.
7 Conclusion, Limitations, and Future Work
Our evaluation of proprietary and open-weight language models in agentic frameworks highlights how underspecificity poses a core challenge in software engineering tasks. Effective performance requires (i) detecting missing information, and (ii) acquiring it through precise, targeted interaction before (iii) attempting a solution with the full information.
Our analysis is subject to a few scope constraints. Underspecificity detection is measured only within the first three turns, as models rarely recover if they fail to engage early. Question quality is approximated via latent vector changes that weigh all information equally, though models may prioritize details differently. Finally, our simulated user proxy may be more cooperative than real users, though we mitigate this by limiting interaction turns and focusing them tightly on the task.
Despite these limitations, several clear trends emerge from our experiments:
  • With a brief round of clarification, leading proprietary models recover much of their fully-specified performance, while earlier open-weight models lag. Recent capable models blur this distinction. However, as models grow more capable, relative gains from interaction diminish, suggesting current training practices inadequately prepare models to dynamically integrate interactive information.
  • LLMs rarely initiate clarification unprompted, and their sensitivity to prompt framing makes them brittle in noisy, real-world contexts.
  • The most effective questions are specific, actionable, and task-level, while vague prompts or implementation details recoverable from the codebase add little value.
Overall, a gap remains between underspecified and fully specified resolution rates. Closing it will require open-weight models to adopt stronger interaction strategies and proprietary models to engage more proactively. As models are trained to perform longer horizon tasks, they must still be trained to appropriately incorporate user inputs into the overall solution. Ambig-SWE provides a blueprint for decomposing resolution into multiple steps, enabling finer-grained analysis of where models succeed or fail. While we focus on software engineering, the methods and insights can extend to other complex, real-world agentic tasks. Thus, our work offers both a diagnostic framework for agent evaluation and a roadmap toward more robust, adaptive, and user-aligned agents that can thrive in underspecified and dynamic environments.
Reproducibility Statement
To ensure the reproducibility of the presented results, this paper provides comprehensive details on the methodology, data generation, and experimental setup. All key components of the proposed framework are described with the intention of enabling replication by an independent research group. The experimental setup is detailed in §2 and full prompts are provided in the Appendix §A. We have also attached the code with the steps to reproduce and the experimental data.
LLM Usage
We used a large language model to assist with polishing the writing style, condensing the content, and improving clarity. All research ideas, methods, experiments, and analyses were developed and conducted by the authors. The LLM did not contribute to scientific content.
Appendix A Appendix
A.1 Agent framework
In this work, we use the OpenHands agent framework for conducting our experiments. OpenHands is a single agent system that has access to tools such as bash terminal, file system, code execution, and browsing (disabled during evaluation). In the SWE-Bench setting, the agent is provided with the issue alongside a detailed prompt which conveys the steps to follow such as exploration, clarification, etc. (as detailed in Appendix A.1.2). Equipped with the above-mentioned tools, the agent interacts with the repository environment inside a Docker container with the required dependencies provided by SWE-Bench. The agent has a maximum number of steps to complete the solution. If finishing early, the agent can call the FinishAction. Upon completion, a git_patch is extracted from the modified files which is later applied to a new environment instance, and the tests associated with the task are run to verify the solution.
A.2 Experimental design
A.2.1 Full setting
A.2.2 Interaction setting
A.2.3 Hidden setting
Metric Mean Median Std Dev
ROUGE-1 Recall 0.179 0.159 0.102
ROUGE-L Recall 0.111 0.094 0.069
Entity Recall 0.085 0.030 0.141
BERTScore F1 -0.111 -0.127 0.194
Table 3: Quantitative comparison of underspecified summaries against full issues using overlap- and semantics-based metrics.
LLM Underspecification Analysis prompt Compare these two texts and identify what INFORMATION is present in the original issue but missing in the problem statement. Focus on factual content differences, not language or writing style differences. Original GitHub Issue:
{original_issue}
Summarized Problem Statement:
{problem_statement}
Instructions: Identify specific pieces of information that appear in the original issue but are absent or underspecified in the problem statement. Focus ONLY on informational content - ignore differences in: Wording or phrasing Writing style or tone Sentence structure Different ways of expressing the SAME information For example: DO include: “Error message ’FileNotFoundError’ is missing” (different information) DO NOT include: “Less detailed explanation of the bug” (same information, different wording) List each missing piece of information as a separate numbered item. Be specific and concrete. Output your analysis as a numbered list within <missing_info></missing_info> tags.
A.3 Statistical methods
A.3.1 Wilcoxon Signed-Rank test
$H_{0}:\tilde{d}\leq 0$
$H_{1}:\tilde{d}>0$
@@ A.3.1 Wilcoxon Signed-Rank test
Qwen 3 Coder Hidden vs Interaction 6.87e-17
Interaction vs Full 5.46e-26
Claude Sonnet 4 Hidden vs Interaction 0.03225
Interaction vs Full 9.0e-29
@@
Table 4: Wilcoxon signed-rank test results for Hidden vs Interaction and Interaction vs Full settings across models.
A.3.2 Compute Requirements
The experiments are conducted using 16 workers in the Remote Runtime (beta) provided in OpenHands which is a cloud-based runtime for parallel execution.
A.4 Naturally underspecified issues
instance_id Issue Discussion
django_django-13952 Migrate signals verbose stdout emissions are not capturable. The migrate command takes a --verbosity flag that is passed down to emit_pre_migrate_signal and emit_post_migrate_signal functions but these are not provided which stdout the output should be directed to. This makes testing migrate -v2 through call_command pollute sys.stdout when it should be directed to the provided stdout as discovered in https://github.com/django/django/pull/13890#pullrequestreview-579320176. Contains concrete technical details (function names, flags), a specific reproducibility scenario (migrate -v2 via call_command), and an external reference link.
sympy__sympy-11794 ASCII printing for Singularity Function. Implementation of ASCII printing for Singularity Functions is needed. Minimal description with no code snippets or reproduction steps, showing that some natural issues are similarly vague despite lacking explicit technical context.
sphinx-doc__sphinx-7234 Support for @singledispatch functions. It would be nice if there was some mechanism to automagically pick up the overloads to a @functools.singledispatch function and list them together. Includes conversational phrasing (stream-of-thought style) and references to a specific Python mechanism, reflecting natural issue-writing patterns.
Table 5: Examples of naturally occurring issues and their characteristic features relevant to underspecification analysis.
A.5 Underspecificity detection prompts
A.6 Question Quality Analysis
$\text{Cosine Distance}(P,Q)=1-\frac{P\cdot Q}{\|P\|\|Q\|}$ (1)
where:
  • $N=1536$ is the dimensionality of the embedding space.
Table 6: Average number of questions asked per interaction.
Model Avg. # Questions
Claude Sonnet 4 4.03
Qwen 3 Coder 6.02